OpenStudy (anonymous):

-10 - (-6) what steps must i follow to solve

OpenStudy (anonymous):

two negatives= a positive, so: -10+6= -4

OpenStudy (anonymous):

so the smallest number is changed positive when it is both negative ?

OpenStudy (anonymous):

Mm no. Sorry I should have given a better explanation :p. A subtraction sign and a negative number change to an addition sign (two negatives). So: 2+2=4 2-2=0 2- (-2)= 2+2=4 2+(-2)= 2-2=0

OpenStudy (anonymous):

ok thank u